If x,y,z are length of the perpandicular dropped from a point inside the triangle of given area A,on the three sides of the triangle then the minimum value of x^2+y^2+z^2 is 4A^2/a^2+b^2+c^2
1
Expert's answer
2019-09-24T11:42:27-0400

Let ABC and H be given triangle and point inside it.

Let "x=HA_1,y=HB_1,z=HC_1" be lengths of the perpendicualars to the sides "a=BC,b=AC,c=AB" respectively. We have "\frac{ax}{2},\frac{by}{2},\frac{cz}{2}" are areas of the triangles "BCH, ACH, ABH" respectively, so "ax+by+cz=2A". We can express "z": "z=\frac{2A-ax-by}{c}".

Then "F(x,y)=x^2+y^2+z^2=x^2+y^2+\bigl(\frac{2A-ax-by}{c}\bigr)^2="

"=\frac{1}{c^2}(x^2c^2+y^2c^2+4A^2+a^2x^2+"

"+b^2y^2-4Aax-4Aby+2abxy)" .

Since "x\ge 0", "y\ge 0" and "z=\frac{2A-ax-by}{c}\ge 0", "F" is defined on the set "U=\{(x,y)\in\mathbb R^2: x\ge0, y\ge0,ax+by\le 2A\}". Since "U" is bounded ("U\subset\bigl[0,\frac{2A}{a}\bigr]\times\bigl[0,\frac{2A}{b}\bigr]" ) and closed, "U" is a compact set by criterion of compactness in "\mathbb R^n" .

Moreover, "F(x,y)=x^2+y^2+z^2\ge 0" for all "x" and "y" . So "F(U)" is bounded from below.

Since "U" is a compact set, "F" is a continuous function bounded from below on "U" , we have that there is the minimum of "F" on "U".

If we want to minimize "F(x,y)" , then we need such "x" and "y" that "\frac{\partial F}{\partial x}(x,y)=0" and "\frac{\partial F}{\partial y}(x,y)=0".

So "0=\frac{\partial F}{\partial x}(x,y)=\frac{2}{c^2}(xc^2+xa^2-2Aa+aby)" and "0=\frac{\partial F}{\partial y}(x,y)=\frac{2}{c^2}(yc^2+yb^2-2Ab+abx)" .

We obtain the linear equation system

"\begin{cases}
xc^2+xa^2-2Aa+aby=0\\
yc^2+yb^2-2Ab+abx=0
\end{cases}"

The system has the only one solution "x_0=\frac{2aA}{a^2+b^2+c^2}" and "y_0=\frac{2bA}{a^2+b^2+c^2}" (so "F(x_0,y_0)" is minimum of "F(x,y)") . Value of "z_0" is "z_0=\frac{2A-ax_0-by_0}{c}=\frac{2cA}{a^2+b^2+c^2}" . So minimum of "F" on "U" is "F(x_0,y_0)=x_0^2+y_0^2+z_0^2=\frac{4A^2}{a^2+b^2+c^2}"